a survey carried out in a town revealed that 25% of the households surveyed owned a laptop computer and 70% owned a desktop computer. in addition, it was found that 12% owned both a laptop and a desktop computer.
if the sample of households from the tow is selected at random determine the proportion that owns Neither a laptop NOR a desktop computer.
1
Expert's answer
2020-03-02T13:54:50-0500

Proportion of households that owns a laptop or a desktop or both:

"P(L\cup D)=P(L)+P(D)-P(L\cap D)=0.25+0.70-0.12=0.83."

Proportion of households that owns Neither a laptop NOR a desktop computer:

"P=1-P(L\cup D)=1-0.83=0.17=17%."%.